OpenStudy (anonymous):

Convert 0.15 g/mL to the unit g/L.

OpenStudy (frostbite):

\[0.15 \frac{ g }{ mL }*10^{3} \frac{ mL }{ L }=150 \frac{ g }{ L }\] I used that: \[1 L=10^{3} mL \rightarrow 1=10^{3} \frac{ mL }{ L }\]